Cobalt-60 has a half-life of approximately 5 years. If there was one microgram present in a sample originally, how many atoms would have decayed after 20 years?

1
Expert's answer
2021-11-21T17:26:19-0500

"N=\frac mMN_A(1-2^{-\frac tT})=9.4\cdot 10^{18}."